2. Suppose demand for lunch box is Qp = 40-P / 5 and the supply of it is Qs = P / 5. Calculate the equilibrium price and quantity of lunch box in this district. Show your answer graphically. (4/20) b. If the price of lunch box cannot be set above%2450. Calculate the "shortage" in this market. Explain your answer with a graph. (2/20) Calculate the price elasticities of demand and supply at quantity demanded and quantity supplied under such a condition.


1
Expert's answer
2020-10-13T18:19:18-0400

The Demands Function is as follows : "Qp = 40 - \frac{p}{5}"


The Supply Function is as follows : "Qs = \frac{p}{5}"


a) In equilibrium,

Qp = Qs

or, "40 - \frac{p}{5}" "=\frac{p}{5}"


or, "40 = \frac{2P}{5}"


or, "P = 40\times \frac{5}{2}"

or, P = 100

and, the corresponding value of Q is 20 units,

b)-) The shortage of the Quantity is given by  

S = Qp - Qs

Now, Qp at P = 24.50 is 35.1 units

and, Qs at P = 24.50 is 4.9 units

hence,  Shortage (S) : 35.1 - 4.9 = 30.2 units

c) Now, Price Elasticity of Demand is the percentage change in Demand due to change in Price

Price Elasticity Of Demand = "(d\frac{Qp}{dp})\times \frac{P}{Q}" at given price and quantity

in this case, "(d\frac{Qp}{dp})" = "-\frac{1}{5}" and so price elasticity in the shortage condition is 0.139 ( putting P=24.5 and Q= 35.1)

Price Elasticity of Supply is the percentage change in Supply due to change in Price

Price Elasticity Of Supply = "(d\frac{Qp}{dp})\times \frac{P}{Q}" at given price and quantity

in this case,  "(d\frac{Qp}{dp})" ="-\frac{1}{5}" and so price elasticity in the shortage condition is 1 ( putting P=24.5 and Q= 4.9)